# A Structural Characterization of the Hit Image in the Motivic Steenrod Algebra

### v2

@@ Abstract.
The motivic hit problem seeks a minimal generating set for the motivic cohomology of classifying spaces as a module over the motivic Steenrod algebra. In this paper, we provide a complete structural classification of the hit image within the top layer of Kameko’s decomposition, revealing a sharp dichotomy governed by parity. Specifically, for degrees $d=k+2d_{1}$ with $d_{1}=(n-1)(2^{k}-1)$, let $V$ denote the span of the monotone translates of Kameko’s monomial $z_{k}$. We prove that the intersection of $V$ with the hit subspace is exactly the even-parity hyperplane. Consequently, the quotient $V/(\text{hits})$ is one-dimensional and is generated by any odd-parity linear combination of these translates. This establishes the ”odd-parity gap” as the precise obstruction to the hit property in this layer. As an arithmetic consequence of this structural result, we identify a new infinite numerical family with $n=2^{r}+1$ and k=n4k=n-4 ($r\geq 5$) where $\beta(d)>n$. This yields a new class of counterexamples to the motivic Peterson-type conjecture, distinct from Kameko’s original family. Furthermore, we show that these structural results and counterexamples persist over any algebraically closed field of characteristic 0.
@@ 1. Introduction
Theorem 1.1 says that $\beta(d)>n$ is a sufficient condition for triviality of the hit quotient. Moreover, the condition is also necessary in the sense that the monomial $x_{1}^{2^{i_{1}}-1}\cdots x_{n}^{2^{i_{n}}-1}$ never lies in $A^{+}(P_{n})$. The interplay between the hit problem and the cohomology of $A$ via Singer’s transfer [7] connects this story to the classical Adams spectral sequence and the stable homotopy groups of spheres; see, for example, Ault [1], Pengelley–Williams [4], the present author [5, 6].
@@
The main objective of this paper is to uncover the sharp structural dichotomy governing the top layer of the motivic hit problem. While previous work focused on the existence of counterexamples, we provide a complete linear-algebraic characterization of the hit image in terms of parity.
@@
    Structural Classification: We isolate a parity functional on the top layer of Kameko’s decomposition and prove that the hit image is exactly the even-parity hyperplane. This implies that the obstruction to the hit problem in this layer is one-dimensional and is completely determined by parity.
@@
    Arithmetic Consequences: We apply this structural insight to the arithmetic of degrees $d=(n-1)(2^{k+1}-2)+k$. Using the equivalence between $\beta(d)>n$ and $\alpha(d+n)>n$ [3, Prop. 2.1], we identify a new infinite family of pairs $(n,k)$ with k=n4k=n-4 (distinct from Kameko’s k=n3k=n-3 case) where the hit quotient is non-zero despite $\beta(d)>n$.
@@
    Generalization: We demonstrate that these results are not specific to $\mathbb{C}$ but persist over any algebraically closed field $K$ of characteristic 0, relying on the stability of the motivic Steenrod algebra under base change.
@@
Theorem 1.4 (Structure of the Top Hit Layer).
For any integers $n\geq 2$ and $1\leq k<n$, let $d=k+2d_{1}$ with $d_{1}=(n-1)(2^{k}-1)$ be the degree of Kameko’s monomial $z_{k}$. Let $M_{1}$ be the set of all translates $\sigma(z_{k})$ under monotone permutations $\sigma\in\mathrm{Mono}(k)$, and let $V$ be their span in the motivic hit quotient $QN_{n}^{d,*}$.
Let $\theta:N_{n}^{d,*}\to V$ be the map extracting the $M_{1}$–component (defined precisely in Section 3). Then there is a linear functional
$\varepsilon:V\to\mathbb{F}_{2}$
such that:
  1. (1)
    Exact Image Characterization: One has
    $\theta\bigl(A^{\sharp}_{+}(N_{n})\cap N_{n}^{d,*}\bigr)=\ker(\varepsilon)\subset V.$
  2. (2)
    Non-hit Criterion: Consequently, any class whose $M_{1}$–component has odd parity is non-hit. The quotient space V/(hit image)V/($\text${hit image}) is exactly one-dimensional.
Remark 1.5 (Minimality of the Obstruction).
Theorem 1.4 shows that the description of the hit image as the even-parity hyperplane is sharp. The “odd-parity gap” is the smallest possible non-trivial quotient, implying that no further reduction by the Steenrod algebra is possible in this layer.
@@ 5.1. Motivic cohomology and the Steenrod algebra over general base fields
with $|\tau|=(0,1)$, $|x_{i}|=(1,1)$ and $|y_{i}|=(2,1)$, and the action of $A_{K}^{*,*}$ is given by the same formulas as over $\mathbb{C}$; see, for example, [8, Sec. 9] together with [2]. In particular, there is a natural isomorphism of bigraded $A^{*,*}$–modules between $M_{n}(K)$ and $M_{n}(\mathbb{C})$, and likewise for their quotients $N_{n}(K)=M_{n}(K)/(\tau)$ and $N_{n}(\mathbb{C})$.

### v3

@@ Abstract.
The motivic hit problem asks for a minimal set of generators of $H^{*,*}(BV_{n};\mathbb{F}_{2})$ as a module over the motivic Steenrod algebra. Masaki Kameko [3] constructed, in a distinguished family of degrees $d=k+2d_{1}$ with $d_{1}=(n-1)(2^{k}-1)$, a decomposition in which the top layer is spanned by the monotone translates of a single monomial $z_{k}$, and he showed that the Bockstein image in this layer is contained in a subspace generated by pairwise sums of these translates. While Kameko obtained a restriction on the possible $Q_{0}$–image in the top layer, we determine the corresponding hit image in $V$ exactly. We make three main contributions as follows:
First, we give an exact structural description of the hit subspace on the top-layer summand $V=\langle\sigma(z_{k})\rangle$ by constructing a parity functional $\varepsilon:V\to\mathbb{F}_{2}$ and proving that $V\cap(\text{hit subspace})=\ker(\varepsilon)$; equivalently, parity yields a complete criterion for hitness in the $M_{1}$–summand and $V/(\text{hits})\cong\mathbb{F}_{2}$. In particular, this computes the quotient $V/(\text{hits})$. We next explain how this parity classification interacts with the numerical condition $\beta(d)>n$. In particular, the counterexamples arise as consequences of the structural theorem.
Second, we combine this classification with the numerical condition $\beta(d)>n$ to obtain a new infinite counterexample family with $n=2^{r}+1$ and k=n4k=n-4 ($r\geq 5$), distinct from Kameko’s k=n3k=n-3 family.
Third, we prove a base-change invariance statement: the parity classification on $V$ (and hence its consequences, including the above family) persists over any algebraically closed field of characteristic 0.
@@ 1. Introduction
Theorem 1.1 says that $\beta(d)>n$ is a sufficient condition for triviality of the hit quotient. Moreover, the condition is also necessary in the sense that the monomial $x_{1}^{2^{i_{1}}-1}\cdots x_{n}^{2^{i_{n}}-1}$ never lies in $A^{+}(P_{n})$. The interplay between the hit problem and the cohomology of $A$ via Singer’s transfer [5] connects this story to the classical Adams spectral sequence and the stable homotopy groups of spheres; see, e.g. [1], [4], [7, 8].
@@
The present paper strengthens Kameko’s analysis of the top layer from an inclusion statement to an exact structural description. In Kameko’s decomposition, the summand
$V=\big\langle\sigma(z_{k})\ \big|\ \sigma\in\mathrm{Mono}(k)\big\rangle$
captures the entire “top layer” contribution in degree $d=k+2d_{1}$. Kameko showed that only certain linear combinations can arise from the Bockstein $Q_{0}$ in this layer, namely combinations generated by pairwise sums $\sigma_{1}(z_{k})+\sigma_{2}(z_{k})$, yielding a containment for the possible $Q_{0}$–image in $V$. What remained open is whether this containment is sharp, i.e. whether it actually exhausts the hit image inside $V$, and hence what the precise obstruction to being hit is. We answer this question by giving a complete parity-based description of the hit image in $V$.
Accordingly, our main contribution in this work is structural: we upgrade Kameko’s top-layer restriction from a containment to an equality by giving an exact parity-based description of the hit image inside $V$. In particular, the counterexample families arise as consequences of this classification, once combined with arithmetic input.
@@
    Parity classification (exact hit image). We construct an explicit linear functional $\varepsilon:V\to\mathbb{F}_{2}$ (the parity) and prove an if-and-only-if criterion: a class in $V$ is hit precisely when it has even parity. Equivalently, the hit image in $V$ is the codimension-one hyperplane $\ker(\varepsilon)$, so that $V/(\text{hits})\cong\mathbb{F}_{2}$.
@@
    Arithmetic application (new counterexample family). As a consequence, once $\beta(d)>n$ holds in these degrees, any odd-parity class in $V$ yields a counterexample to the motivic Peterson-type conjecture. Using the equivalence between $\beta(d)>n$ and $\alpha(d+n)>n$ [3, Prop. 2.1], we establish a new infinite family with k=n4k=n-4 (distinct from Kameko’s k=n3k=n-3 case) for which $\beta(d)>n$.
@@
    Base change. We show that the parity classification on $V$ is stable under extension of algebraically closed base fields of characteristic 0, and hence all of its consequences persist under base change.
@@
Theorem 1.4 (Parity classification in the top layer).
For integers $n\geq 2$ and $1\leq k<n$, let $d=k+2d_{1}$ with $d_{1}=(n-1)(2^{k}-1)$, and let $V$ be the $\mathbb{F}_{2}$–span in $QN_{n}^{d,*}$ of the monotone translates $\sigma(z_{k})$ with $\sigma\in\mathrm{Mono}(k)$. Let $\theta:N_{n}^{d,*}\to V$ be the projection onto the $M_{1}$–component (defined in Section 3).
There exists a nonzero linear functional $\varepsilon:V\to\mathbb{F}_{2}$ such that:
  1. (i)
    Exact hit image in $V$:
  2. (ii)
    If and only if criterion: For $v\in V$, one has $v$ is hit in $QN_{n}^{d,*}$ if and only if $\varepsilon(v)=0$. In particular $V/(\text{hits})\cong\mathbb{F}_{2}$, generated by any $v\in V$ with $\varepsilon(v)=1$.
Remark 1.5 (Sharpness).
Theorem 1.4 upgrades Kameko’s restriction [3] on the possible $Q_{0}$–image in the top layer to an equality statement: the hit image inside $V$ is a single explicit codimension-one hyperplane. Thus the obstruction in this summand is minimal (one-dimensional) and is completely detected by parity. Equivalently, the theorem computes the one-dimensional quotient $V/(\text{hits})$, and provides an effective parity criterion: a class in $V$ is hit if and only if it has even parity.
@@ 5.1. Motivic cohomology and the Steenrod algebra over general base fields
with $|\tau|=(0,1)$, $|x_{i}|=(1,1)$ and $|y_{i}|=(2,1)$, and the action of $A_{K}^{*,*}$ is given by the same formulas as over $\mathbb{C}$; see, for example, [6, Sec. 9] together with [2]. In particular, there is a natural isomorphism of bigraded $A^{*,*}$–modules between $M_{n}(K)$ and $M_{n}(\mathbb{C})$, and likewise for their quotients $N_{n}(K)=M_{n}(K)/(\tau)$ and $N_{n}(\mathbb{C})$.